OpenStudy (anonymous):

Find the intercepts for the graph of the equation given. 4x-7y=84. Find the x-intercept. Find the y-intercept. I keep working on this problem, but I keep getting stuck, I don't know what I am doing wrong.

OpenStudy (ghazi):

divide whole equation by 84 you'll get 4/84*x-7/84*y=1 and then you'll have x/21-y/12=1 hence the denominator of x and y are the intercepts..because the intercept form of equation of line is x/a+y/b=1 where a and b are intercepts

OpenStudy (ghazi):

you'll have (21,-12)

OpenStudy (anonymous):

Thanks, I forgot all about dividing. Now I just have to check and see if it simplifies.

OpenStudy (ghazi):

okay...hope you'll get the answer

OpenStudy (anonymous):

I hope so to.

OpenStudy (ghazi):

what you got?

OpenStudy (anonymous):

Well, to be honest right now I feel very dumb lol. I forgot how to simplify these things.

OpenStudy (ghazi):

noo worries it's just 4/84 for x intercept and 7/84 for y now after solving you'll get x intercept = 21 and y= -12 for further query check out above mentioned comments

OpenStudy (anonymous):

Lol, alright. I hate math like this it always gives me a headache.

OpenStudy (ghazi):

it's easy ...play with it :)

OpenStudy (anonymous):

Lol, not easy for me. I can never get it to stick in my head.

OpenStudy (ghazi):

okay give a try ..and if you're stuck...just ask for it... :)

OpenStudy (anonymous):

Thanks.

OpenStudy (ghazi):

YW